OpenStudy (anonymous):

simplify bt reducing the radical 12√125x^3

OpenStudy (anonymous):

is 12 the index or a coefficient?

OpenStudy (anonymous):

\[\sqrt[12]{125x^3}\] or \[12\sqrt{125x^3}\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

i will guess index, and then write \[\sqrt[12]{(5x)^3}\] 3 goes in to 12 4 times so answer is \[\sqrt[4]{5x}\]

OpenStudy (anonymous):

clear what i did yes? divide 12 by 3 to change the index from 12 to 4

OpenStudy (anonymous):

yes